OpenStudy (anonymous):

Natasha bought 20 liters of soft drinks for her party. What would be an equivalent amount in quarts? 1 quart (qt) .................. 0.95 liter (L) Round off your decimal answer to 2 decimal places. Enter only significant zeros. would the answer be 0.05qrts?

OpenStudy (anonymous):

think of setting up this question like this, you know that 0.95/1 gives you the relationship between litters and quarts, and you are trying to figure out how many quarts 20 litters is. so if 0.95litres = 1 quart; 20litres = how many quart? So now you know that 0.95/1 = 20/something lets call it x. Because the ratio of liters to quarts must be maintained when forming an equation. So if 0.95 = 20/x ;;; 0.95x = 20. So what does x equal?